OpenStudy (anonymous):

please help explain this to me. solve for the x's by factoring x(3x+1)=2 I know I'm suppose to subtract the two to both side to make the equation equal zero but not sure whats next

OpenStudy (kropot72):

Expand x(3x + 1) then subtract 2 from each side. Do you recognise what you then have?

OpenStudy (anonymous):

4x-1=0?

OpenStudy (anonymous):

never mind i see my mistake

OpenStudy (anonymous):

4X+X-2=0 ?

OpenStudy (kropot72):

\[x(3x+1)=3x ^{2}+x\]

OpenStudy (anonymous):

huh I dont know why i keep going to 4x in stead of 3xsquared

OpenStudy (anonymous):

how did you do the little squared thinggy

OpenStudy (kropot72):

I used the Equation button below the answer box. \[3x \times x=3x ^{2}\]

OpenStudy (anonymous):

i see my mistake now, thank so much, i gave you a medal

OpenStudy (kropot72):

You're welcome :). You should reach the following: \[3x ^{2}+x -2=0\] Canyou factorise this to find the two solutions for x?

OpenStudy (anonymous):

yes 3x-2=0 is x=two thirds and x+1=0 is x=-1

OpenStudy (kropot72):

Well done!!

OpenStudy (anonymous):

ok now can you help me with finding the square root of something.?

OpenStudy (kropot72):

Yes.Bring it on!